\begin{thebibliography}{199}
\addcontentsline{toc}{section}{References}

\bibitem{milan}  This paper is an updated version of E.H.~Lieb:   \textit{The
Stability of Matter and Quantum Electrodynamics}, Milan Jour. Math.
{\bf 71}, 199-217 (2003).  An earlier version appears in the proceedings
of the Werner Heisenberg Centennial, December, 2001 in Munich. Springer
Verlag (in press).


\bibitem{ahs} J. Avron, I. Herbst, B. Simon:  \textit{Schr\"odinger
operators with magnetic fields III}, Commun. Math. Phys. \textbf{79},
529- 572 (1981).

\bibitem{bfg} L. Bugliaro, J. Fr\"ohlich, G. M. Graf:  \textit{Stability
of quantum electrodynamics with nonrelativistic matter},
Phys. Rev. Lett. \textbf{77}, 3494-3497 (1996).

\bibitem{bfs} V. Bach, J. Fr\"ohlich, I. M. Sigal: \textit{Spectral
analysis for systems of atoms and molecules coupled to the quantized
radiation field}. Comm.  Math. Phys. \textbf{207} 249-290 (1999).

\bibitem{conlon} J. G. Conlon:  \textit{The ground state of a classical
gas}, Commun. Math. Phys. \textbf{94},  439-458 (1984).


\bibitem{cly} J. G. Conlon, E. H. Lieb, H.-T. Yau: \textit{The $N^{7/5}$
law for charged bosons}, Commun. Math. Phys. \textbf{116}, 417-448 (1988).

\bibitem{daubechies} I. Daubechies: \textit{An uncertainty
principle for fermions with generalized kinetic energy},
Commun. Math. Phys. \textbf{90}, 511-520 (1983).

\bibitem{daubechieslieb} I. Daubechies,E. H. Lieb: \textit{One
electron relativistic molecules with Coulomb interaction},
Commun. Math. Phys. \textbf{90}, 497-510 (1983).

\bibitem{dyson1} F. J. Dyson: \textit{Ground state energy of a finite
system of charged particles}, J. Math. Phys. \textbf{8}, 1538-1545 (1967).

\bibitem{dl} F. J. Dyson, A. Lenard: \textit{Stability of matter I and
    II}, J. Math. Phys. \textbf{8}, 423-434 (1967), \textbf{9},
  1538-1545 (1968).

\bibitem{eps} W. D. Evans, P. P. Perry, H. Siedentop: \textit{The spectrum
of relativistic one-electron atoms according to Bethe and Salpeter},
Commun. Math. Phys. \textbf{178}, 733-746 (1996).

\bibitem{feff} C. Fefferman: \textit{Stability of Coulomb systems in a
magnetic field}, Proc. Nat. Acad. Sci. USA, \textbf{92}, 5006-5007 (1995).

\bibitem{fl} C. Fefferman, R. de la Llave: \textit{Relativistic stability
of matter. I}. Rev. Mat. Iberoamericana \textbf{2}, 119-213 (1986).

\bibitem{fll} J. Fr\"ohlich, E. H . Lieb, M. Loss: \textit{Stability
of Coulomb systems with magnetic fields I.  The one-electron atom},
Commun. Math. Phys. \textbf{104}, 251-270 (1986). See also E. H. Lieb,
M. Loss: \textit{Stability of Coulomb systems with magnetic Fields
II.  The many-electron atom and the one-electron molecule}, Commun.
Math. Phys. \textbf{104}, 271-282 (1986).

\bibitem{gll}  M. Griesemer, E. H. Lieb, M. Loss: \textit{Ground states
in non-relativistic quantum electrodynamics}, Invent. Math. \textbf{145},
557-595 (2001).

\bibitem{gt} M. Griesemer, C. Tix: \textit{Instability of
pseudo-relativistic model of matter with self-generated magnetic field},
J. Math. Phys.  \textbf{40}, 1780-1791 (1999).

\bibitem{hainzsei} C. Hainzl and R. Seiringer: \textit{Mass
renormalization and energy level shifts in non-relativistic QED},
Adv. Theor. Math. Phys.  \textbf{6}, 847-871 (2002).

\bibitem{hs} G. Hoever, H. Siedentop: \textit{The Brown-Ravenhall
operator}, Math. Phys. Electronic Jour. \textbf{5}, no. 6 (1999).

\bibitem{lieb1}  E. H. Lieb: \textit{The stability of matter},
Rev. Mod. Phys.  \textbf{48},  553-569 (1976).

\bibitem{lieb3} E. H. Lieb:  \textit{The $N^{5/3}$ law for bosons},
Phys. Lett. \textbf{70A}, 71-73  (1979).

\bibitem{lieb2} E. H. Lieb:  \textit{The stability of matter:  From
atoms to stars},
 Bull. Amer. Math. Soc. \textbf{22}, 1-49 (1990).


\bibitem{lieblebowitz} E. H. Lieb, J.L. Lebowitz: \textit{The
existence of thermodynamics for real matter with Coulomb forces},
Phys. Rev. Lett. \textbf{22}, 631-634 (1969).

\bibitem{anal}  E. H. Lieb, M.~Loss: \textit{Analysis}, American
Mathematical Society (1997).

\bibitem{liebloss} E. H. Lieb, M. Loss: \textit{Stability of a Model of
Relativistic Quantum Electrodynamics}, Commun. Math. Phys. \textbf{228},
561-588 (2002).  arXiv math-ph/0109002, mp\_arc 01-315.

\bibitem{LLmass} E. H. Lieb, M. Loss: \textit{A Bound on Binding
Energies and Mass Renormalization in Models of Quantum Electrodynamics},
J. Stat. Phys.  \textbf{108}, 1057-1069 (2002).  arXiv math-ph/0110027.

\bibitem{LLatoms} E. H. Lieb, M. Loss: \textit{Existence of
Atoms and Molecules in Non-relativistic Quantum Electrodynamics},
Adv. Theor. Math. Phys. (in press).  arXiv math-ph/0307046.

\bibitem{lls} E. H. Lieb, M.~Loss, H.~Siedentop: \textit{Stability
of relativistic matter via Thomas-Fermi theory}, Helv.~Phys.~Acta
\textbf{69}, 974-984 (1996).

\bibitem{llsolovej} E. H. Lieb, M. Loss, J. P. Solovej: \textit{Stability
of matter in magnetic fields}, Phys.~Rev.~Lett. \textbf{75}, 985-989
(1995).

\bibitem{lss} E. H. Lieb, H.~Siedentop, J.~P.~Solovej: \textit{Stability
and instability of relativistic electrons in magnetic fields},
J.~Stat.~Phys. \textbf{89}, 37-59 (1997). See also \textit{Stability of
relativistic matter with magnetic fields}, Phys. Rev. Lett.  \textbf{79},
1785-1788 (1997).

\bibitem{liebsimon}  E. H. Lieb, B. Simon: \textit{Thomas-Fermi theory
revisited}, Phys. Rev. Lett. \textbf{31}, 681-683 (1973).

\bibitem{liebthirring}  E. H. Lieb, W. Thirring:  \textit{Bound for
the kinetic energy of fermions which proves the stability of matter},
Phys. Rev. Lett. \textbf{35}, 687-689 (1975). Errata \textbf{35}, 1116
(1975).

\bibitem{liebyau}  E. H. Lieb, H.-T. Yau: \textit{The stability and
instability of relativistic matter}, Commun. Math. Phys. \textbf{118},
177-213 (1988).  See also \textit{Many-body stability implies a bound
on the fine structure constant}, Phys. Rev. Lett. \textbf{61}, 1695-1697
(1988).


\bibitem{ly} M. Loss, H.-T. Yau:  \textit{Stability of Coulomb systems
with magnetic fields III. Zero energy bound states of the Pauli operator},
Commun.  Math. Phys. \textbf{104}, 283-290 (1986).

\bibitem{planck} M. Planck: \textit{Zur Theorie des Gesetzes der
Energieverteilung im Normalspektrum}, Verhandlung der Deutschen
Physikalischen Gesellschaft \textbf{2}, 237-245 (1900).

\bibitem{ZG} G. Zhislin: \textit{A study of the spectrum of the
Schr\"odinger operator for a system of several particles}, Trudy Moscov
Mat. Obsc.  \textbf{9},  81-120 (1960).




\end{thebibliography}

\documentclass[12pt]{article}

\newcommand{\uDelta}{\Delta}
\newcommand{\uPi}{\mathrm{\Pi}}
\newcommand{\uLambda}{{\Lambda}}
\newcommand{\A}{\mathbf{A}}
\newcommand{\B}{\mathbf{B}}
\newcommand{\x}{\mathbf{x}}
\newcommand{\bk}{\mathbf{k}}
\newcommand{\R}{\mathbf{R}}

\newcommand{\s}{\boldsymbol{\sigma}}
\newcommand{\p}{\mathbf{p}}
\newcommand{\da}{\boldsymbol{\alpha}}
\newcommand{\D}{\mathrm{d}}
\newcommand{\eps}{\boldsymbol{\varepsilon}}

\usepackage{amsthm,amsfonts,amsmath}

\begin{document}

\title{\bf Quantum Mechanics, The Stability of Matter and 
Quantum Electrodynamics}
\author{\vspace{5pt} Elliott H. Lieb \\
\vspace{-4pt} 
\normalsize Departments of Mathematics and Physics,
Princeton University \\
\normalsize P.O. Box 708, Princeton, NJ 08544-0708, USA\\
\normalsize lieb@princeton.edu \\ } 

\date{\small December 30, 2003}

\maketitle             

\renewcommand{\thefootnote}{}

\renewcommand{\thefootnote}{}
\footnotetext{This paper is an extended version of a talk at the DMV
Jahrestagung in Rostock, 17 November, 2003 \cite{milan}.}
\renewcommand{\thefootnote}{}
\footnotetext{
\copyright\, 2003 by the author. This article may be reproduced, in its
entirety, for non-commercial purposes.}
\renewcommand{\thefootnote}{}
\footnotetext{Work partially
supported by U.S. National Science Foundation
grant PHY 0139984-A01.}

\begin{abstract}
Much progress has been made in the last few decades in developing the
necessary mathematics for understanding the full implications of the
quantum-mechanical many-body problem and why the material world appears
to be as stable as it is despite the serious $-1/|x|$ singularity of the
Coulomb potential that attracts negative electrons to positive atomic
nuclei. Many problems remain, however, especially the understanding of
the interaction of matter and the quantized radiation field discovered
by Planck in 1900.  A short review of some of the main topics is given.
\end{abstract}

\section{Introduction}

This paper is a brief survey of the quantum-mechanical many-body
problem, especially the question of the interaction of matter with
radiation.  The quantum-mechanical revolution of the 1920's brought
with it many successes, but also a few problems that have yet to be
resolved. The realization that there were a few problems with the
simple textbook theory surfaced three or four decades ago. Since then
some of the mathematical questions have been answered, but some big
ones remain.  This brief overview might, it is hoped, encourage
some mathematicians to look into this fascinating topic.

We begin with a sketch of the topics that will concern us
here.

\subsection{Triumph  of Quantum Mechanics}\label{triumph}

One of the basic problems of classical physics (after the discovery of
the point electron by Thomson and of the (essentially) point nucleus
by Rutherford) was the stability of atoms. Why do the electrons in an
atom not fall into the nucleus? Quantum mechanics explained this fact.
It starts with the classical Hamiltonian of the system (nonrelativistic
kinetic energy for the electrons plus Coulomb's law of electrostatic
energy among the charged particles). By virtue of the non-commutativity
of the kinetic and potential energies in quantum mechanics the stability
of an atom -- in the sense of a finite lower bound to the energy -- was a
consequence of the fact that any attempt to make the electrostatic energy
very negative would require the localization of an electron close to the
nucleus and this, in turn, would result in an even greater, positive,
kinetic energy.

Thus, the basic stability problem for an atom was solved by an
inequality that says that  $\langle 1/|x| \rangle$, 
the expected value of $1/|x|$, can be made large
only at the expense of making the kinetic energy, which is
proportional to $\langle p^2 \rangle$, even larger. A fundamental
hypothesis of quantum mechanics is that $p$ is represented by the
differential operator $-i \hbar \nabla$ with $\hbar =h/2\pi$ and
$h=$Planck's constant.  In elementary presentations of the subject it
is often said that the mathematical inequality that ensures this fact
is the famous uncertainty principle of Heisenberg (proved by Weyl),
which states that $\langle p^2 \rangle
\langle x^2 \rangle
\geq (9/8)\hbar^2 $.

While this principle is mathematically rigorous     it is actually
insufficient for the purpose, as explained, e.g., in
\cite{lieb1,lieb2}, and thus gives only a heuristic explanation of the
power of quantum mechanics to prevent collapse.  A more powerful
inequality, such as Sobolev's inequality (\ref{sobolev}), is needed
(see, e.g., \cite{anal}). The utility of
the latter is made possible by Schr\"odinger's representation of
quantum mechanics (which earlier was a somewhat abstract theory of
operators on a Hilbert space) as a theory of differential operators on
the space of square integrable functions on $\mathbb{R}^3$. The
importance of Schr\"odinger's representation is sometimes
underestimated by formalists, but it is of crucial importance because
it permits the use of functional analytic methods, especially
inequalities such as Sobolev's, which are not easily visible on the
Hilbert space level. These methods are essential for the developments
reported here.

To summarize, the understanding of the stability of atoms and ordinary
matter requires a formulation of quantum mechanics with two ingredients:
\noindent \begin{itemize} \item 
A Hamiltonian formulation in order to
have a clear notion of a lowest possible (ground state) energy. 
Lagrangian formulations,
while popular, do not always lend themselves to the identification of
that quintessential quantum mechanical notion of a ground state energy.
In quantum mechanics a Hamiltonian is not a function on phase space but rather
a (pseudo-) differential operator.
\item 
A formulation in terms of concrete function spaces instead of
abstract Hilbert spaces so that the power of mathematical analysis can
be fully exploited.  
\end{itemize}

\subsection{Some Basic Definitions}

As usual, we shall denote the lowest energy (eigenvalue) of a quantum
mechanical system by $E_0$. (More generally, $E_0$ denotes the infimum
of the spectrum of the Hamiltonian $H$ in case this infimum is not an
eigenvalue of $H$ or is $-\infty$.) Our intention is to investigate
arbitrarily large systems, not just atoms. In general we suppose that
the system is composed of $N$ electrons and $K$ nuclei of various kinds.
Of course we could include other kinds of particles but $N$ and $K$
will suffice here. $N=1$ for a hydrogen atom and $N=10^{23}$ for a mole of
hydrogen.  We shall use the following terminology for two notions of stability:
\begin{eqnarray} \label{firststab}
&E_0&
> -\infty  \quad\qquad\qquad\qquad \mathrm{Stability\ of\  the\  first\
kind,}\\ 
&E_0& > C(N+K)  \qquad\qquad \mathrm{Stability\
of\  the\  second\  kind} \label{secondstab}
\end{eqnarray} 
for some constant $C\leq 0$ that is independent of $N$ and $K$, but
which may depend on the physical parameters of the system (such as the
electron charge and mass). Usually, $C<0$, which means that there is a
positive binding energy per particle.

Stability of the second kind is absolutely essential if quantum
mechanics is going to reproduce some of the basic features of the
ordinary material world: The energy of ordinary matter is extensive
(i.e., it is proportional to the number of particles),
the thermodynamic limit exists (i.e., the $N\to \infty$ limit exists)
and the
laws of thermodynamics hold. Bringing two stones together might
produce a spark, but not an explosion with a release of energy
comparable to the energy in each stone.  Stability of the second kind
does not guarantee the existence of the thermodynamic limit for the
free energy, but it is an essential ingredient \cite{lieblebowitz}
\cite[Sect. V]{lieb1}. 

It turns out that stability of the second kind cannot be taken for
granted, as Dyson discovered \cite{dyson1}. If Coulomb forces are
involved, then {\it the Pauli exclusion principle is essential.}
(This means that the $L^2$ functions of $N$ variables, $\Psi (x_1,\>
x_2, \dots, x_N)$, \ $x_i\in {\mathbb{R}^3}$, is antisymmetric under all
transpositions $x_i \leftrightarrow x_j$. Particles, like electrons, whose
functions $\Psi$ obey this principle are called {\it fermions}. Particles
whose $\Psi $ functions are symmetric under permutations are called
{\it bosons}.)

Charged bosons are {\it not stable} because for them $E_0\sim -N^{7/5}$
(nonrelativistically) and $E_0 = -\infty$ for large, but finite $N$
(relativistically, see Sect.  \ref{relmanybody}). While  positively
charged bosons exist in the form of atomic nuclei, negatively charged,
long-lived bosons do not exist in nature. This is a good thing in view of
the instability just mentioned.

\subsection{The Electromagnetic Field}

A second big problem handed down from classical physics was
the `electromagnetic mass' of the electron. This poor creature has
to drag around an infinite amount of electromagnetic energy that
Maxwell burdened it with. Moreover, the electromagnetic field itself
is quantized -- indeed, that fact alone started the whole revolution
\cite{planck}.

While quantum mechanics accounted for stability with Coulomb forces
and Schr\"odinger led us to think seriously about the `wave function
of the universe', physicists shied away from talking about the wave
function of the particles in the universe {\it and} the
electromagnetic field in the universe. It is noteworthy that
physicists are happy to discuss the quantum mechanical many-body
problem with external electromagnetic fields non-perturbatively, but
this is rarely done with the quantized field. The quantized field
cannot be avoided because it is needed for a correct description of
atomic radiation, the laser, etc. However, the interaction of matter
with the quantized field is almost always treated perturbatively or
else in the context of highly simplified models (e.g., with two-level
atoms for lasers).

The quantized electromagnetic field greatly complicates the stability
of matter question. It requires, ultimately, mass and charge
renormalizations.  At present such a complete theory does not exist,
but a theory {\it must} exist because matter exists and because we
have strong experimental evidence about the manner in which the
electromagnetic field interacts with matter, i.e., we know the
essential features of a  Hamiltonian that adequately accounts for the 
low energy processes that exist in every day life.  In short, nature
tells us that it must be possible to formulate a self-consistent
quantum electrodynamics (QED) {\it non-perturbatively,} (perhaps with
an ultraviolet, or high frequency, cutoff of the field at a few MeV).  
It should not be
necessary to have recourse to quantum chromodynamics (QCD) or some
other high energy theory to explain ordinary matter.  

Physics and other natural sciences are successful because physical
phenomena associated with each range of energy and other parameters
are explainable to a good, if not perfect, accuracy by an appropriate
self-consistent theory. This is true whether it be hydrodynamics,
celestial dynamics, statistical mechanics, etc.  If low energy physics
(atomic and condensed matter physics) is not explainable by a
self-consistent, non-perturbative theory on its own level one can
speak of an epistemological crisis.

Some readers might say that QED is in good shape. After all, it
accurately predicts the outcome of some very high precision experiments
(Lamb shift, $g$-factor of the electron). But the theory does not
really work well when faced with the problem, which is explored here,
of understanding the many-body ($N\approx 10^{23}$) problem and the
stable low energy world in which we spend our everyday lives.

\subsection{Relativistic Mechanics}

When the classical kinetic energy of a particle,
$p^2/2m$, is replaced by its relativistic
version $\sqrt{p^2c^2 +m^2c^4} $ the stability question becomes
much more complicated, as will be seen later. It turns out that even
stability of the first kind is not easy to obtain and it depends on the
values of the physical constants, notably the fine structure constant
\begin{equation} \label{alpha}
\alpha= e^2/\hbar c =1/137.04 \ ,
\end{equation} 
where $-e$ is the
electric charge of the electron.

For ordinary matter relativistic effects are not dominant but they are
noticeable. In large atoms these effects severely change the innermost
electrons and this has a noticeable effect on the overall electron
density profile.  Therefore, some version of relativistic mechanics is
needed, which means, presumably, that we must know how to replace
$p^2/2m$ by the Dirac operator (see (\ref{dirac})).

The combination of relativistic mechanics plus the electromagnetic field
(in addition to the Coulomb interaction) makes the stability problem
difficult and uncertain.  Major aspects of this problem have been worked
out in the last few years (about 35) and that is the subject of this paper.

\section{Nonrelativistic Matter without the Magnetic Field}
\label{nomagnet}

Maxwell's equations define the electric and magnetic fields in terms of
potentials. While the equations determine the fields, the potentials
are not determined uniquely; the choice of potentials is called the
choice of gauge.  We work in the `Coulomb' gauge for the electromagnetic
field. Despite the assertion that quantum mechanics and quantum field
theory are gauge invariant, it seems to be essential to use this gauge,
even though its relativistic covariance is not as transparent as that
of the Lorentz gauge. The reason is the following.

The Coulomb gauge is the gauge in which electrostatic part of the
interaction of matter with the electromagnetic field is just the
conventional Coulomb ``action at a distance'' potential $V_c$ given by
(\ref{coul}) below (in energy units $mc^2$ and length units the
Compton wavelength $\hbar/mc$). This part of the interaction depends
only on the coordinates of the particles and not on their velocities.
The dependence of the interaction on velocities, or currents, comes
about through the magnetic part of the interaction.  Despite
appearances, this picture is fully Lorentz invariant (even if it is
not gauge invariant).
\begin{eqnarray} \label{coul}
V_c = - \sum_{i=1}^N \sum_{k=1}^K \frac{Z_k}{|\x_i - \R_k|}
+\sum_{1\leq i < j \leq N}\frac{1}{ |\x_i-\x_j|} 
+\sum_{1\leq k < l \leq K}\frac{Z_kZ_l}{|\R_k-\R_l|} \  .
\end{eqnarray}
The first sum is the interaction of the electrons (with dynamical
coordinates $\x_i$) and fixed nuclei located at $\R_k$ of positive
charge $Z_k$ times the (negative) electron charge $e$. The second is
the electron-electron repulsion and the third is the nucleus-nucleus
repulsion. The nuclei are fixed because they are so massive
relative to the electron that their motion is irrelevant.  It could be
included, however, but it would change nothing essential.  Likewise, there
is no nuclear structure factor because if it were essential for stability
then the size of atoms would be the size of nuclei, about $10^{-13}$
cm, instead of about $10^{-8}$ cm, contrary to what is observed.

Although the nuclei are fixed points the constant $C$ in the stability of
matter (\ref{secondstab}) is required to be independent of the $\R_k$'s.
Likewise (\ref{firststab}) requires that $E_0$ have a finite lower bound
that is independent of the $\R_k$'s.

For simplicity of exposition we shall assume here that all the 
$Z_k$ are identical, i.e., $Z_k=Z$.

The magnetic field, which will be introduced later, is described by a 
vector potential $\A(x)$ which is a dynamical variable in the Coulomb gauge.
The magnetic field is $\B=\mathrm{curl}\A$.

There is a basic physical distinction between electric and magnetic forces
which does not seem to be well known, but which motivates this choice of
gauge.  In electrostatics ``like charges repel'' while in magnetostatics
``like currents attract''. A consequence of these facts is that the
correct magnetostatic interaction energy can be obtained by minimizing the
energy functional $\frac{1}{2} \int B^2 - \int \mathbf{j}\cdot \A$ with
respect to the vector field $\A$, where $\mathbf{j}$ is the 
electric current density.  The positive electrostatic energy, on the other
hand, {\it cannot} be obtained by a minimization principle with respect
to the field (e.g., minimizing $\frac{1}{2} \int | \boldsymbol{\nabla}
\phi|^2 -\int \phi \varrho$ with respect to $\phi$).

The Coulomb gauge, which puts in the electrostatics correctly, by
hand, so to speak, and allows us to minimize the total energy with respect
to the $\A$ field, is the gauge that gives us the correct physics and
is consistent with the ``quintessential quantum mechanical notion of
a ground state energy'' mentioned in Sect. \ref{triumph}. In any other
gauge one would have to look for a critical point of a Hamiltonian rather
than a true global minimum.

The type of Hamiltonian that we wish to consider in this
section is
\begin{equation}
H_N= T_N+ \alpha V_c\ . 
\end{equation}
Here, $T_N$ is the kinetic energy of the $N$ electrons and has the form
\begin{equation}\label{kinetic}
T_N= \sum_{i=1}^N T_i \ ,
\end{equation}
where $T_i$ acts on the coordinate of the $i^{th}$ electron.
The nonrelativistic choice is $T=p^2$ with $\p=-i\boldsymbol{\nabla}$
and $p^2 = -\uDelta$ in appropriate units.

\subsection{Nonrelativistic Stability for Fermions}

The problem of stability of the second kind for nonrelativistic
quantum mechanics was recognized in the early days by a few
physicists, e.g., Onsager, but not by many. It was not solved until
1967 in one of the most beautiful papers in mathematical physics by
Dyson and Lenard \cite{dl}. 

They found that the Pauli principle, i.e., Fermi-Dirac statistics, is
essential.  Mathematically, this means that the Hilbert space is the
subspace of antisymmetric functions, i.e.,
${\mathcal{H}}^{\mathrm{phys}} =
\wedge^N  L^2({\mathbb{R}}^3; {\mathbb{C}}^2)$. This is how the 
Pauli principle is interpreted post-Schr\"odinger; Pauli invented his
principle a year earlier, however!

Their value for $C$ in (\ref{secondstab}) was rather high, about
$-10^{15}$ eV (electron volts) for $Z=1$. (The ground state energy of
a hydrogen atom is -13 eV.) The situation was improved later by Thirring
and myself \cite{liebthirring} to about $-20$ eV for $Z=1$ by introducing
an inequality that holds only for the kinetic energy of
fermions (not bosons) in an arbitrary state $\Psi$.

\begin{equation}\label{lt}
\langle \Psi, T_N \Psi \rangle \geq (const.) \int_{\mathbb{R}^3}
\varrho_\Psi(\x)^{5/3} \, \D^3 \x \  ,
\end{equation}
where $\varrho_\Psi$ is the one-body density in the (normalized)
fermionic wave function $\Psi$ (of space and spin) given by an
integration over $(N-1)$ coordinates and $N$ spins as follows.
\begin{equation}\label{rho}
\varrho_\Psi(\x) = N\sum_{\sigma_1,\dots ,
\sigma_N} \int_{\mathbb{R}^{3(N-1)}}|\Psi(\x,\, \x_2,...,\x_N; 
\sigma_1,\dots \sigma_N)|^2\, \D^3\x_2\cdots \D^3\x_N  \  .
\end{equation} 

Inequality (\ref{lt}) allows one simply to reduce the quantum mechanical
stability problem to the stability of Thomas-Fermi theory, which was worked
out earlier by Simon and myself \cite{liebsimon}. 

The older inequality of Sobolev, mentioned in Sect. \ref{triumph},
\begin{equation}\label{sobolev}
\langle \Psi, T_N \Psi \rangle \geq (const.)  \left(\int_{\mathbb{R}^3}
\varrho_\Psi(\x)^{3} \, \D^3 \x \right)^{1/3}\ ,
\end{equation}
is not as useful as (\ref{lt}) for the many-body problem because its
right side is proportional to $N$ instead of $N^{5/3}$. It is, however,
strong enough to yield the stability of a system, like an atom, that
has only a few electrons.

It is amazing that from the birth of quantum mechanics until  1967 none
of the luminaries of physics had quantified the fact that electrostatics
plus the uncertainty principle {\it do not suffice} for stability of the
second kind, and thereby make thermodynamics possible (although they
do suffice for the first kind). See Sect. \ref{bose}.  It was noted,
however, that the Pauli principle was responsible for the large sizes
of atoms and bulk matter (see, e.g., \cite{dyson1,dl}).

\subsection{Nonrelativistic Instability for Bosons}\label{bose}

What goes wrong if we have charged bosons instead of fermions?
Stability of the first kind (\ref{firststab}) holds in the
nonrelativistic case, but (\ref{secondstab}) fails. If we assume the nuclei 
are infinitely massive, as before, and $N=KZ$ then $E_0 \sim -N^{5/3}$
\cite{dl,lieb3}. To remedy the situation we can let the nuclei have finite
mass (e.g., the same mass as the negative particles). Then, as Dyson
showed \cite{dyson1}, $E_0 \leq -(const.)N^{7/5}$. This calculation
was highly non-trivial! Dyson had to construct a variational function
with pairing of the Bogolubov type in a rigorous fashion and this took
several pages.

Thus, finite nuclear mass improves the situation, but not enough. The
question whether $N^{7/5}$ is the correct power law remained open for
many years.  A lower bound of this type was needed and that was finally
obtained in \cite{cly}.

The results of this Section \ref{nomagnet} can be summarized by saying
that stability of the hydrogen atom is one thing but stability of
many-body physics is something else \thinspace !

\section{Relativistic Kinematics (no magnetic field)}\label{rel}

The next step is to try to get some idea of the effects of relativistic
kinematics, which means replacing $p^2$ by $\sqrt{p^2+ 1}$ in non-quantum
physics. (Recall that $mc^2=1$ in our units.) The simplest way to do this
is to substitute $\sqrt{p^2+ 1}$ for $T$ in (\ref{kinetic}).  The Dirac
operator will be discussed later on, but for now this choice of $T$
will suffice. Actually, it was Dirac's choice before he discovered his
operator and it works well in some cases. For example, Chandrasekhar used
it successfully, and accurately, to calculate the collapse of white dwarfs
(and later, neutron stars).

Since we are interested only in stability, we may, and shall, substitute
$|\p| = \sqrt{-\uDelta}$ for $T$. The error thus introduced is bounded
by a constant times $N$ since $|\p|<\sqrt{p^2+ 1}< |\p|+1$ (as an
operator inequality).  Our Hamiltonian is now $H_N =\sum_{i=1}^N |\p_i|
+\alpha V_c$.

\subsection{One-Electron Atom}\label{oneelectron}

The touchstone of quantum mechanics is the Hamiltonian for `hydrogen'
which is, in our case,
\begin{equation}\label{relhyd}
H= |\p| -Z\alpha/|\x| = \sqrt{-\uDelta} -Z\alpha/|\x| \ .
\end{equation}

It is well known (also to Dirac) that the analogous operator with
$|\p|$ replaced by the Dirac operator (\ref{dirac}) 
ceases to make sense when
$Z\alpha >1$.  Something similar happens for (\ref{relhyd}).
\begin{equation} \label{crit}
E_0= 
\begin{cases} 
0 &\text{if $Z\alpha \leq 2/\pi$;} \\
-\infty   &\text{if $Z\alpha >2/\pi$ .}  
\end{cases}
\end{equation}

The reason for this behavior is that both $|\p|$ and
$|\x|^{-1}$ scale in the same way. Either the first term in
(\ref{relhyd}) wins or the second does.

A result similar to (\ref{crit}) was obtained in \cite{eps}
for the free Dirac operator $D(0)$ in place of $|\p|$, but with 
the wave function
$\Psi$ restricted to lie in the positive spectral subspace of $D(0)$. 
Here, the critical value is $\alpha Z \leq (4\pi)/(4+ \pi^2) >2/\pi$.

The moral to be drawn from this is that relativistic kinematics plus
quantum mechanics is a `critical' theory (in the mathematical sense).
This fact will plague any relativistic theory of electrons and the
electromagnetic field -- primitive or sophisticated.

\subsection{Many Electrons and Nuclei}\label{relmanybody}

When there are many electrons is it true that the condition
$Z\alpha \leq const.$ is the only one that has to be considered?
The answer is no! One {\it also} needs the condition that 
$\alpha $ itself must be small, regardless of how small $Z$ might be.
This fact can be called a `discovery' but actually it is an overdue
realization of some basic physical ideas. It should have been
realized shortly after Dirac's theory in 1927, but it does  not
seem to have been noted until 1983 \cite{daubechieslieb}. 

The underlying physical heuristics is the following. With $\alpha $
fixed, suppose $ Z\alpha = 10^{-6}\ll 1$, so that an atom is stable,
but suppose that we have $2\times 10^{6}$ such nuclei. By bringing
them together at a common point we will have a nucleus 
with $ Z\alpha =2$ and one electron suffices to cause collapse into it.
Then (\ref{firststab}) fails. What prevents this from happening,
presumably, is the nucleus-nucleus repulsion energy which goes to 
$+\infty$ as the nuclei come together. 
But this repulsion energy is proportional to $(Z\alpha)^2/\alpha$ and, 
therefore, if we regard $Z\alpha$ as fixed we see that $1/\alpha $
must be large enough in order to prevent collapse.

Whether or not the reader believes this argument, the mathematical
fact is that there is a fixed, finite number $\alpha_c \leq 2.72 $ 
 (\cite{liebyau}) so that when $\alpha > \alpha_c$
(\ref{firststab}) fails for {\it every} positive $Z$ and for 
every $N\geq 1$ (with or without the Pauli principle). 

The open question was whether (\ref{secondstab}) holds for {\it all}
$N$ and $K$ if $Z\alpha $ and $\alpha $ are both small enough. The
breakthrough was due to Conlon \cite{conlon} who proved
(\ref{secondstab}), for fermions, if $Z=1$ and $\alpha <
10^{-200}$. The situation was improved by Fefferman and de la Llave
\cite{fl} to $Z=1$ and $\alpha < 0.16$.  Finally, the expected correct
condition $Z\alpha
\leq 2/\pi$ and $\alpha < 1/94$ was obtained in \cite{liebyau}. (This
paper contains a detailed history up to 1988.) The situation was
further improved in \cite{lls}. The multi-particle version of the use
of the free Dirac operator, as in Sect. \ref{oneelectron}, was treated 
in \cite{hs}.

Finally, it has to be noted that charged bosons are {\it always} unstable
of the first kind (not merely the second kind, as in the nonrelativistic
case) for {\it every} choice of $Z>0, \alpha > 0$. E.g., 
there is instability if  $Z^{2/3}\alpha N^{1/3} > 36$ (\cite {liebyau}). 

We are indeed fortunate that there
are no stable,  negatively charged  bosons.  

\section{Interaction of Matter with Classical Magnetic Fields}\label{magfields}

The magnetic field $\B$ is defined by a vector potential $\A(\x)$ and 
$\B(\x) =\mathrm{curl}\, \A(\x)$. In this section we take a first step 
(warmup exercise) by
regarding $\A$ as classical, but indeterminate, and we introduce the
classical field energy 
\begin{equation}\label{classfield}
H_f = \frac{1}{8\pi}\int_{\mathbb{R}^3} B(\x)^2 \D x \ .
\end{equation}

 The Hamiltonian is now
\begin{equation}\label{fieldham}
H_N(\A) = T_N(\A)+ \alpha V_c + H_f \ ,
\end{equation}
in which the kinetic energy operator has the 
form (\ref{kinetic}) but depends on $\A$.
We now define $E_0$ to be  the infimum of 
$\langle \Psi,\  H_N(\A) \Psi \rangle$ both with respect to
$\Psi $ {\it and with respect to} $\A$. 

\subsection{Nonrelativistic Matter with Magnetic Field}

The simplest situation is merely `minimal coupling' without spin, namely,
\begin{equation}
T(\A) = |\p +\sqrt\alpha \A(\x)|^2
\end{equation}
This choice does not change any of our previous results qualitatively.
The field energy is not needed for stability.  On the one-particle
level, we have the `diamagnetic inequality' $\langle \phi,\ |\p+\A(\x)
|^2 \phi \rangle \geq \langle |\phi|,\ p ^2 |\phi| \rangle$. The same
holds for $|\p+\A(\x)|$ and $|\p|$. More importantly, inequality
(\ref{lt}) for fermions continues to hold (with the same constant)
with $T(\A)$ in place of $p^2$. (There is an inequality similar to
(\ref{lt}) for $|\p|$, with $5/3$ replaced by $4/3$, which also
continues to hold with minimal substitution \cite{daubechies}.)

The situation gets much more interesting if spin is included. This takes
us a bit closer to the relativistic case. The kinetic energy operator
is the Pauli operator
\begin{equation} \label{pauli-fierz}
T^P(\A) = |\p + \sqrt\alpha\; \A(\x)|^2 + \sqrt\alpha \;\B(\x)\cdot \s\ ,
\end{equation}
where $\s$ is the vector of $2\times 2$ Pauli spin matrices and
$L^2({\mathbb{R}^3}) $ is replaced by $L^2({\mathbb{R}^3}; {\mathbb{C}^3}) $

\subsubsection{One-Electron Atom}

The stability problem with $T^P(\A)$ is complicated, even for a
one-electron atom.  Without the field energy $H_f$ the Hamiltonian is
unbounded below.  (For fixed $\A$ it is bounded but the energy tends
to $-\infty$ like $-(\log B)^2$ for a homogeneous field
\cite{ahs}.) The field energy saves the day, but the result is surprising 
\cite{fll}
(recall that we must minimize the energy with respect to $\Psi$ {\it and}
$\A$):
\begin{equation}
|\p + \sqrt\alpha\; \A(\x)|^2 + \sqrt\alpha \;\B(\x)\cdot \s
-Z\alpha/|\x| +H_f
\end{equation}
{\it is bounded below if and only if $Z\alpha^2 \leq C$,} where $C$ is some 
constant that can be bounded as $1<C <9\pi^2/8$. 

The proof of instability \cite{ly} is difficult and requires the
construction of a zero mode (soliton) for the Pauli operator, i.e., a
finite energy magnetic field and a {\it square integrable} $\psi$ such
that 
\begin{equation} \label{zeromode} 
T^P(\A)\psi =0\ .  
\end{equation}
The usual kinetic energy
$|\p+\A(\x)|^2$ has no such zero mode for any $\A$, even when 0 is
the bottom of its spectrum.

The original magnetic field \cite{ly} that did the job in
(\ref{zeromode}) is independently interesting, geometrically (many
others have been found since then). 
$$ \B(\x) = \frac{12 }{(1 + |\x|^2)^3}
[(1-x^2) \mathbf{w} + 2(\mathbf{w}\cdot \x) \x + 2 \mathbf{w}\land \x]
$$ 
with $\vert \mathbf{w} \vert = 1$. The field lines of this magnetic
field form a family of curves, which, when stereographically projected
onto the 3-dimensional unit sphere, become the great circles in what
what is known as the Hopf fibration.

Thus, we begin to see that nonrelativistic matter with magnetic fields
behaves like relativistic matter without fields -- to some extent.

The moral of this story is that a magnetic field, which we might think
of as possibly self-generated, can cause an electron to fall into the
nucleus. The uncertainty principle cannot prevent this, not even for
an atom!

\subsubsection{Many Electrons and Many Nuclei}

In analogy with the relativistic (no magnetic field) case, we can see that
stability of the first kind fails if $Z\alpha^2$ {\it or} $\alpha$ is 
too large. The heuristic reasoning is the same and the proof is similar.

We can also hope that stability of the second kind holds
 if both $Z\alpha^2$ {\it and} $\alpha$ are small enough.
The problem is complicated by the fact that it is the field energy $H_f$ that
will prevent collapse, but there there is only one field energy while there
are $N\gg 1$ electrons.

The hope was finally realized, however. Fefferman \cite{feff} proved stability
of the second kind for  $H_N(\A)$ with the Pauli $T^P(\A)$ for
$Z=1$ and ``$\alpha$ sufficiently small''. A few months later it
was proved \cite{llsolovej} for $Z\alpha^2 \leq 0.04$ and $\alpha \leq
0.06$.  With $\alpha =1/137$ this amounts to $Z\leq 1050$. This very
large $Z$ region of stability is comforting because it means that
perturbation theory (in $\A$) can be reliably used for this particular
problem.

Using the results in \cite{llsolovej}, Bugliaro, Fr\"ohlich and Graf
\cite{bfg} proved stability of the same nonrelativistic 
Hamiltonian -- but with an ultraviolet cutoff, quantized magnetic
field whose field energy is described below. (Note: No cutoffs are
needed for classical fields.)

There is also the very important work of Bach, Fr\"ohlich, and Sigal
\cite{bfs} who showed that this nonrelativistic Hamiltonian with
ultraviolet cutoff, quantized field {\it and} with sufficiently small
values of the parameters has other properties that one expects. E.g.,
the excited states of atoms dissolve into resonances and only the ground
state is stable. The infrared singularity notwithstanding, the ground
state actually exists (the bottom of the spectrum is an eigenvalue);
this was shown in \cite{bfs} for small parameters and in \cite{gll},
\cite{LLatoms}
for all values of the parameters. (See Sect. \ref{atoms}.)

\section{Relativity Plus Magnetic Fields} \label{relmag}

As a next step in our efforts to understand QED and the many-body problem
we introduce relativity theory along with the classical magnetic field.

\subsection{Relativity Plus Classical Magnetic Fields} \label{relmagc}

Originally, Dirac and others thought of replacing $T^P(\A)$ by
$\sqrt{T^P(\A) +1} $ but this was not successful mathematically and does not
seem to conform to experiment. Consequently, we introduce the Dirac operator
for $T$ in (\ref{kinetic}), (\ref{fieldham})
\begin{equation}\label{dirac}
D(\A) = \da \cdot \p + \sqrt{\alpha}\ \da \cdot \A(\x)
 + \beta m \ ,
\end{equation}
where $\da $ and $\beta$ denote the $4\times 4$ Dirac matrices
and  $ \sqrt{\alpha}$ is the electron charge as before. (This notation
of $\da$ and $\alpha$ is historical and is not mine.) 
The Hilbert space for $N$ electrons is now changed to
\begin{equation}\label{oldh}
\mathcal{H} = \wedge^N L^2(\mathbb{R}^3; \mathbb{C}^4)\  .
\end{equation}

The well known problem with $D(\A) $ is that it is unbounded below, and so
we cannot hope to have stability of the first kind, even with $Z=0$.
Let us imitate QED (but without pair production or renormalization) by
restricting the electron wave function to lie in the positive 
spectral subspace of a Dirac operator.

Which Dirac operator?

There are two natural operators in the problem. One is $D(0)$, the
free Dirac operator. The other is $D(\A)$ that is used in the
Hamiltonian. In almost all formulations of QED the electron is defined
by the positive spectral subspace of $D(0)$. Thus, we can define
\begin{equation} \label{hphys}
\mathcal{H}^{\mathrm{phys}} = P^+\ \mathcal{H} = \uPi_{i=1}^N  
\pi_i  \, \mathcal{H} \  ,
\end{equation}
where $ P^+=\uPi_{i=1}^N \pi_i $, and $\pi_i$ is the projector
of onto the positive
spectral subspace of $D_i(0) = \da \cdot \p_i 
+ \beta m$, the free Dirac operator for the $i^{\mathrm{th}}$ electron.
We then restrict the allowed wave functions in the variational principle
to those $\Psi$ satisfying
\begin{equation}
\Psi = P^+\ \Psi  \qquad\quad i.e., \ \Psi \in  
\mathcal{H}^{\mathrm{phys}}   \  .
\end{equation}

Another way to say this is  that we replace the Hamiltonian
(\ref{fieldham}) by $P^+ \, H_N \, P^+$ on $\mathcal{H}$ and look 
for the bottom of its spectrum.

It turns out that this prescription leads to disaster! While the use
of $D(0)$ makes sense for an atom, it fails miserably for the
many-fermion  problem, as discovered in \cite{lss} and refined in
\cite{gt}. The result is: 

{\it For all $\alpha >0$ in (\ref{dirac}) (with or without the Coulomb
term $\alpha V_c$) one can find $N$ large enough so that $E_0=
-\infty$.}

In other words, the term $\sqrt{\alpha}\, \da\cdot \A$ in the Dirac operator
can cause an instability that the field energy cannot prevent.

It turns out, however, that the situation is saved if
one uses the positive spectral subspace of the Dirac operator $D(\A)$ 
to define an electron. (This makes the concept  of an electron $\A$ 
dependent, but when we make the vector potential into a dynamical quantity
in the next section, this will be less peculiar since there will be no
definite vector potential but only a fluctuating quantity.) 
The definition of the physical Hilbert space is as in (\ref{hphys}) but with
$\pi_i$ being the projector onto the positive subspace of the 
full Dirac operator
$D_i(\A) = \da \cdot \p_i + \sqrt{\alpha}\ \da \cdot \A(\x_i)
 + \beta m $. Note that these $\pi_i$ projectors commute with each other
and hence their product $P^+$ is a projector.

The result \cite{lss} for this model ((\ref{fieldham}) with the Dirac
operator and the restriction to the positive spectral subspace of $D(\A)$)
is reminiscent of the situations we have encountered
before:

{\it If $\alpha $ and $Z$ are small enough stability of the second kind
holds for this model.} 

Typical stability values that are rigorously established \cite{lss} are
$Z\leq 56$ with $\alpha =1/137$ or $\alpha \leq 1/8.2$ with $Z=1$. 

\section{Quantized Electromagnetic Fields}

Let us now try to analyze some of the problems connected with the
quantization of the electromagnetic field. The great discovery of Max
Planck \cite{planck}, which was the first step in the new quantum theory,
was that the energy of the electromagnetic field came in quantized units.
The energy unit of electromagnetic waves of frequency $\nu$ is $h\nu$,
and in terms of wave number $k$ (i.e., the wave is proportional to $\exp
(ik\cdot x)$) it is $\hbar c |k|$ since $2\pi \nu / |k| = c =$ speed
of light.

We begin with the problem of generalizing the results in the 
previous subsection to the quantized field.

\subsection{Relativity Plus Quantized Magnetic Field}

The obvious next step is to try to imitate the strategy of Sect.
\ref{relmagc} but with the quantized $\A$ field. This was done
in \cite{liebloss}. The quantized $A$ field is described by an 
operator-valued Fourier transform as
\begin{equation}\label{apot}
\A(\x) = \frac{1}{2\pi} \sum_{\lambda=1}^2 \int_{|\bk|\leq \uLambda}
\frac{\eps_\lambda(\bk)}{\sqrt{|\bk|}} \Big[
a_\lambda(\bk) e^{i\bk\cdot \x} + a_\lambda^{\ast}(\bk) e^{-i\bk\cdot 
\x}\Big]
\D^3 \bk \  ,
\end{equation}
where $\uLambda$ is the ultraviolet cutoff on the wave-numbers $|\bk|$.
The operators $a_{\lambda}, a^{\ast}_{\lambda}$
satisfy the usual canonical commutation relations
\begin{equation}
[a_{\lambda}(\bk), a^{\ast}_{\nu} (\mathbf{q})] = \delta ( \bk-\mathbf{q})
\delta_{\lambda, \nu}\ , ~~~ [a_{\lambda}(\bk), a_{\nu} (\mathbf{q})] =
0, \quad {\mathrm{etc}}
\end{equation}
and the vectors $\eps_{\lambda}(\bk)$ are two
orthonormal polarization vectors perpendicular to $\bk$ and to each other.

The field energy $H_f$ is now given by a normal-ordered version of 
(\ref{classfield})
\begin{equation}\label{eq:fielden}
H_f = \sum_{\lambda=1,2} ~ \int_{\mathbb{R}^3} ~ |\bk|\ 
a_\lambda^{\ast}(\bk)a_\lambda(\bk) \D^3 \bk
\end{equation}

The Dirac operator is the same as before, (\ref{dirac}). Note that 
$D_i(\A)$ and $D_j(\A)$ still commute with each other (since
$\A(\x)$ commutes with $\A(\mathbf{y})$). This is important because it allows
us to imitate Sect. \ref{relmagc}.

In analogy with (\ref{oldh}) we define 
\begin{equation}
\mathcal{H} = \wedge^N L^2(\mathbb{R}^3; \mathbb{C}^4)
\otimes \mathcal{F}\ , 
\end{equation}
where $\mathcal{F}$ is the Fock space for the
photon field. We can then define the {\it physical} Hilbert space as before
\begin{equation}
\mathcal{H}^{\mathrm{phys}} = \Pi\ \mathcal{H} = \uPi_{i=1}^N  
\pi_i  \, \mathcal{H}\  ,
\end{equation}
where the projectors $\pi_i$ project onto the 
positive spectral subspace of either $D_i(0)$ or
$D_i(\A)$.

Perhaps not surprisingly, the former case leads to catastrophe, 
as before.
This is so, even with the ultraviolet cutoff, which we did not have
in Sect.  \ref{relmagc}. Because of the cutoff the catastrophe is milder
and involves instability of the second kind instead of the first kind.
This result relies on a coherent state construction in \cite{gt}.

The latter case (use of $D(\A)$ to define an electron) leads to stability
of the second kind if $Z$ and $\alpha $ are not too  large. Otherwise,
there is instability of the first kind.  The rigorous estimates are
comparable  to the ones in Sect. \ref{relmagc}.

Clearly, many things have yet to be done to understand the stability of
matter in the context of QED. Renormalization and pair production have
to be included, for example.

The results of this section suggest, however, that a significant change
in the Hilbert space structure of QED might be necessary.  We see that
it does not seem possible to keep to the current view that the Hilbert
space is a simple tensor product of a space for the electrons and a Fock
space for the photons. That leads to instability for many particles (or
large charge, if the idea of `particle' is unacceptable). The `bare'
electron is not really a good physical concept and one must think of
the electron as always accompanied by its electromagnetic field. Matter
and the photon field are inextricably linked in the Hilbert space
$\mathcal{H}^{\mathrm{phys}} $.


\bigskip
\bigskip

The following tables \cite{liebloss} summarize some of the results of
this and the previous sections \bigskip \bigskip

\centerline{\bf Electrons defined by projection onto the positive}
\centerline{\bf subspace of $D(0)$, the free Dirac operator}

\bigskip
\begin{tabular}{l||c|c|}
&Classical or quantized field  & Classical or quantized field \\
 &\quad without cutoff $\uLambda$ &  with cutoff $\uLambda$ \\
& $\alpha >0$ but arbitrarily small. & $\alpha >0$ but arbitrarily small.\\
 & & \\
\hline\hline
Without Coulomb&  Instability of &  Instability  of \\
potential $\alpha V_c$  & the first kind& the second kind \\
 \hline
With Coulomb &  Instability of &  Instability of \\
potential $\alpha V_c$ & the first kind &  the second kind \\
 \hline\hline
\end{tabular}

\bigskip\bigskip

\smallskip

\centerline{\bf Electrons defined by projection onto the  positive}
\centerline{\bf subspace of $D(\A)$, the Dirac operator with field}

\bigskip
\quad\quad 
\begin{tabular}{l||c|c|}
&\multicolumn{2}{c| } {Classical field with or without cutoff $\uLambda$ } \\
&\multicolumn{2}{c| } {or quantized field with  cutoff $\uLambda$} \\
   &\multicolumn{2}{c | }   {}  \\
\hline\hline
Without Coulomb& \multicolumn{2}{c |} {The Hamiltonian is positive}  \\
potential $\alpha V_c$ &  \multicolumn{2}{c |}  {}     \\
 \hline
&\multicolumn{2}{c |} {Instability of the first kind
when either} \\
With Coulomb & \multicolumn{2}{c|} {$\alpha$ or $Z\alpha$ is too large}\\
\cline{2-3}
potential $\alpha V_c$ & \multicolumn{2}{c|}{Stability of the second kind
when}\\
& \multicolumn{2}{c| }{both $\alpha$ and $Z\alpha$ are small enough}\\
\hline \hline
\end{tabular}


\subsection{Mass Renormalization}

\smallskip
In both classical and quantum electrodynamics there is a problem
of mass renormalization. This means that when a charge is accelerated its
accompanying electromagnetic field is also accelerated and acts like an
additional mass. The `bare mass' of the particle (which is the mass that 
appears in the Hamiltonian) must be chosen so that the
final, physical mass (as measured in experiments) 
agrees with the physically measured value. 

For a point particle, the additional mass is infinity, classically. For QED 
it is also infinite, but the divergence is less rapid as the radius 
of the charge goes to zero. In any case, with a finite ultraviolet
cutoff $\Lambda$ the additional mass is finite, but it is 
far from clear that,
for each $\Lambda > 0$ one can adjust the bare mass (while keeping  it
positive) to give the correct physical mass. Opinions differ on this point
and very little is known rigorously about the problem outside of
perturbation theory. See \cite{hainzsei}.

There are two ways to define  mass renormalization. Take one particle
($N=1$) and then \underbar{\tt either}

1. Find the bottom of the spectrum of $T+H_f$ under the condition that
the total momentum of particle plus field 
is $p$. Call it $E(p)$ and write, for small $p$,
$$
E(p)=E(p=0) + {p^2 / 2m_{\rm physical}}
$$

\underbar{{\tt or else}}
\smallskip
                                                                              
2. Compute the binding energy of hydrogen ($N=1, K=1, Z=1$).
Call it $E_0$  and set
$$
E_0 = {m_{\rm physical} c^2 \alpha^2 / 2\hbar^2}
$$
                                                                         
The first way is the usual one; the second is motivated by the earliest
experiment in quantum mechanics.  These two definitions are not 
the same. In any case, we \cite{LLmass} can now obtain non-trivial bounds
on the binding energy (in the context of the Schr\"odinger Hamiltonian
or the Pauli Hamiltonian interacting with the quantized field)
and thereby get some bounds on the renormalized
mass using definition 2. For large cutoff $\Lambda$, these bounds
differ in their $\Lambda$ dependence from what might be expected from
perturbation theory.
                                                                   
\section{Existence of Atoms in Non-relativistic QED}\label{atoms}

One of the most recent topics concerns the seemingly trivial question
of the existence of atoms. In some sense this question is the opposite
of the stability of matter question.

The Hamiltonian we shall use to describe an atom or
molecule   with $N$ electrons is 
\begin{equation} \label{atomham}
H_N = \sum_{i=1}^N T^P_i(A) +\alpha V_c +H_f
\end{equation}
where $ T^P_i(A)$ is the Pauli kinetic energy operator
(\ref{pauli-fierz}),
but $A$ is the quantized magnetic field given by (\ref{apot}), and
$H_f$ is the energy of the quantized field given by (\ref{eq:fielden}).
As before, $V_c$ is the Coulomb potential (\ref{coul}) of some
fixed nuclei whose total nuclear charge is denoted by $Z=\sum Z_j$.

To show the existence of stable atoms we need to establish two
things about $H_N$

\smallskip

1.  The ground state energy (bottom of the spectrum)
of $H_N$ is lower than that of  $H_{N'}$ i.e., of 
a system with $N'<N$ electrons (with the remaining $N-N'$
electrons being allowed to escape to infinity). This is called 
\textit{the binding condition}.

\smallskip

2. The bottom of the spectrum of $H_N$ is actually an
eigenvalue, i.e., Schr\"odinger's equation  has a square
integrable solution with $E=$ the bottom of the spectrum.

In the case of the Schr\"odinger equation without the
field, problem 1. was solved by Zhislin in 1960 for the 
case $N<Z+1$, which
includes the neutral molecule. He did this by using a localization
technique, whose positive localization energy ($r^{-2}$) is
more than offset by the Coulomb attraction ($-r^{-1}$)  of a positively
charged system
($Z-N'$) to a negatively charged electron. The existence
of the ground state (problem 2.) follows from standard 
arguments because in this case
the bottom of the spectrum is negative while the bottom of the 
essential spectrum (which, in this case, is the bottom of the continuum)
starts at zero. Thus, there is a gap 
in the spectrum
and the technique of taking weak limits easily yields a non-zero
eigenfunction \cite{anal}. 

When we turn on the interaction with the quantized magnetic field the
situation changes significantly. One major difference is that the
bottom of the essential spectrum is now the bottom of the spectrum 
because we
can always create photons with arbitrarily small energy (recall that
the energy of a photon with momentum $k$ is $|k|$). Therefore, if a
ground state exists it necessarily  lies at the bottom of the essential
spectrum and is not isolated. Eigenvalues in the continuum are notoriously
difficult to handle, even for the simple Schr\"odinger operator.

A second major difference is that it is necessary to localize the $A$
field as well as the electrons. This localization costs an energy
$r^{-1}$, not $r^{-2}$ as before, essentially because the field energy
is proportional to $|k|$ instead of $k^2$. Thus, the field
localization competes with the Coulomb attraction.

Problems 1. and 2. were solved in \cite{bfs} under the condition that
$\alpha$ and $\Lambda$ are small enough. 

The first general result, valid for all values of the various
constants, was in \cite{gll}, where it was shown that 2. holds whenever
1. holds. 

Finally, 1. was shown to hold for all values of the 
constants \cite{LLatoms}
under the same natural condition as Zhislin's, i.e., $N<Z+1$.

\end{document}